OpenStudy (anonymous):

A system of equations is shown below: x + 3y = 5 (equation 1) 7x – 8y = 6 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 3x + y = 5 and 8x –7y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 8x – 5y = 11 and 7x – 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 15x + 13y = 17 and 7x – 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations –13x + 15y = 17 and 7x – 8y = 6 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@sammixboo @abb0t @omg1help @Princer_Jones @LiteNing1337 @KendrickLamar2014 @k_lynn @love10129151 @priecetheveil

OpenStudy (anonymous):

@LyricalDevonne @liaba @sleepyhead314

OpenStudy (anonymous):

maths is not mine subject infact i hate it @therealsamalbert

OpenStudy (anonymous):

@iamabarbiegirl @invinceble @Pamela16 @gibbs @quickstudent

OpenStudy (gibbs):

idk sowwy